# PhoniTale: Phonologically Grounded Mnemonic Generation for Typologically Distant Language Pairs

### v2

@@ Abstract
Vocabulary acquisition poses a significant challenge for second-language (L2) learners, especially when learning typologically distant languages such as English and Korean, where phonological and structural mismatches complicate vocabulary learning. Recently, large language models (LLMs) have been used to generate keyword mnemonics by leveraging similar keywords from a learner’s first language (L1) to aid in acquiring L2 vocabulary. However, most of this research has focused on native English speakers learning other languages, rather than the reverse. In this paper, we present PhoniTale, a novel cross-lingual mnemonic generation system that retrieves L1 keyword sequence based on phonological similarity and uses LLMs to generate mnemonics. We evaluate PhoniTale using both automated metrics and human evaluations, comparing its output to mnemonics created by humans and by previous automated approaches. To assess practical effectiveness, we also conduct a short-term recall test measuring mnemonic helpfulness. Our findings show that PhoniTale performs comparably to human-authored mnemonics. We also highlight key areas for future improvement in mnemonic quality and methodology.
@@ 1 Introduction
Typically, the L1 phrase corresponding to a given L2 term to be memorized is manually designed, however this is a laborious process that scales poorly, necessitating the development of automated mechanisms to compose these phrases. Methods for automated generation of such keywords began with TransPhoner, which leverages the International Phonetic Alphabet (IPA) International Phonetic Association (1949) and hand-crafted heuristics to retrieve pronunciation-similar L1 keyword for target English words, resulting in significant recall gains Savva et al. (2014).
Leveraging these methods, recent studies employ large language models (LLMs) for automated keyword generation. SmartPhone fed the TransPhoner keyword into GPT-3 to automatically generate verbal cues and DALL·E to generate visual cues Lee and Lan (2023). Lee et al. (2024) introduced an overgenerate-and-rank approach, where LLMs overgenerate keyword sequences and verbal cues, and then rank them according to multiple different criteria. Balepur et al. (2024) aligned mnemonics with user preferences by fine-tuning Llama 2 for personalization and cost-efficiency.
These prior work focus predominantly on Indo-European L1-L2 language pairs that have substantial phonological overlap Savva et al. (2014); Lee and Lan (2023). However, typologically distant language pairs, such as English-Korean, present unique challenges that remain underexplored. English and Korean exhibit four major phonological mismatches that makes mnemonic generation challenging. First, the orthographic systems differ in dimensionality because English prints letters linearly, while Korean arranges jamo into two-dimensional syllable blocks Park and Li (2009). Second, Korean forbids consonant clusters within a syllable, so epenthetic vowels must be inserted when adapting cluster-rich English words, which expands the syllable count Kang (2003); Kenstowicz (2005). Third, certain English phonemes such as /θ/ have no direct Korean counterpart and are usually replaced with /s/ or /t/ Tak (2012); Kim and Kochetov (2011). Fourth, the two languages exhibit phonemic contrast differences: Korean employs a three-way lenis, fortis, and aspirated stop contrast, whereas English distinguishes only voiced versus voiceless stops, and treats aspiration as a position-dependent allophone Kang (2014); Kang et al. (2022). These fundamental differences complicate the generation of phonologically faithful keyword sequence when Korean speakers learn English.
@@
In this paper, we introduce a mnemonic generation system for language learning, PhoniTale. Our approach employs a greedy search through phonetically and syllabically approximated L2 sequences to identify the most suitable L1 keyword sequence. Specifically, we first transliterate L2 phonemes into L1-adapted sequences, segment these into syllables according to L1 phonological constraints, and then select keywords that maximize phonetic similarity while preserving syllabic structure. Unlike previous approaches that rely heavily on LLMs for keyword generation, we utilize LLMs only for verbal cue generation, while our specialized modules handle the cross-lingual phonological alignment. This design addresses the unique challenges posed by typologically distant languages, improves scalability, and mitigates the risk of hallucinations. Through systematic evaluation including both automated metrics and human studies with short-term recall tests, we demonstrate that PhoniTale achieves comparable performance to human-authored mnemonics.
@@ 2 Problem Statement
PhoniTale performs the task of retrieving cross-lingual phonologically similar keyword sequence and using them to construct a verbal cue for a given target word, following the process illustrated in Figure 1. Let be a word in the L2, and let denote its meaning. The goal is to retrieve a L1 keyword sequence that are phonologically similar to segments of , and to use to construct a verbal cue , where is the space of natural-language expressions in L1.
@@ 3.1.2 Segmentation
The second stage combines these syllables into at most two segments for lengthy words, addressing the fundamental challenge that one-to-one mapping with proves ineffective due to phoneme combinations or syllable structures in L2 that are absent in L1 Daland and Zuraw (2013). This binary constraint prevents excessive fragmentation while preserving phonetic similarity and phonological coherence. The resulting segment sequence serves as input to the subsequent keyword sequence retrieval module, facilitating phonologically informed matching between and .
@@ 3.2 Verbal Cue Generation
Second, we discard the Age-of-Acquisition (AoA) ranking criterion from Lee et al. (2024) as it does not generalize effectively to cross-lingual contexts. The AoA of a word in L2 fails to reliably reflect its familiarity in L1. We retain only the context completeness criterion, calculating this by masking the target word in the verbal cue and prompting GPT-4o to generate five probable candidates. We then compute the average cosine similarity between FastText Bojanowski et al. (2016) embeddings of these candidates and the target word, trained on Korean corpus data Lee (2020). This approach quantifies how effectively the verbal cue provides context for learning the target word’s meaning.
@@ 4.1 Dataset
We construct a training dataset comprising 2,870 English-Korean word pairs with aligned IPA transcriptions. English vocabulary items originate from standardized GRE preparation materials, including official Educational Testing Service guides and commercial preparation resources Magoosh (2021); Princeton Review (2020). Our dataset features English words, their Korean transliteration equivalents, English IPA transcriptions, and Korean IPA transcriptions. We extract Korean transliterations from the AHA Dictionary Aha Dictionary (2022) and convert these transliterations to Korean IPA using the rule-based hangul_to_ipa tool Nam (2021), which implements standard Korean phonological processes Shin et al. (2012). Finally, we verify and segment the Korean IPA representations according to syllabic boundaries, creating aligned syllable-level annotations between the English and Korean IPA forms.
For test dataset, we use the book KSS Gyeong (2020) as our baseline for human-authored verbal cues, which are designed for native Korean speakers learning English. The vocabulary in this book targets advanced-level standardized tests, including government employee entrance exams, university transfer admissions, the TOEFL Educational Testing Service , and the TEPS Seoul National University . From this vocabulary, we construct a test set of 36 words sampled from the book.
@@ 4.2 Cross-lingual IPA Transliteration
To understand how our model addresses the phonological divergence between English and Korean, we analyze attention patterns (See Appendix Figure 6). The visualizations reveal the model’s strategies for managing cross-linguistic challenges: English affricates decompose into multiple Korean consonants; compatible sounds maintain one-to-one mappings; English diphthongs expand to accommodate Korean’s vowel inventory; and syllable structures adapt to conform to Korean phonotactic constraints. These patterns confirm the model’s ability to dynamically adjust its mapping strategy based on input characteristics.
@@ 4.3 Segmentation
Our model achieves a perfect boundary-level F1 score of 1.00 when compared to these manually annotated boundaries on the system’s transliteration outputs. This result indicates exceptional precision in identifying syllable junctures within the predicted L1-adapted phoneme sequence.
@@ 5.1 Dataset
We use same dataset mentioned in 4.1 for evaluation. We replicate Lee et al. (2024) in a same way for generating cross-lingual verbal cues, except that we translate prompts that were originally written in English with in-context examples for English-to-English learning into Korean, and use in-context examples from KSS to compare with PhoniTale. Hereafter, we refer to the human-authored book as KSS, Lee et al. (2024) as OGR, and PhoniTale as PHT.
@@ 5.2.1 Keyword Sequence
We evaluate quality on the following three aspects: phonetic similarity, keyword omission, and keyword modification. We evaluate phonetic similarity using our IPA-based contrastive model, which measures how closely the concatenated Korean keyword sequence resemble sound of the English target word.
@@ 5.3.1 Keyword Sequence
KSS, authored by human, one possible reason for the low phonetic similarity stems from incorporating L2 morphological elements such as prefixes (e.g., “in-”, “re-”) and suffixes (e.g., “-cracy”) instead of using L1 keywords. Although these elements appear in the keyword sequence, L1 learners need to interpret them in Korean during memorization. For instance, even when re (meaning “again”) is included as part of a keyword, learners must mentally translate it into the Korean equivalent of “again” during the association process. The added step of translation increases cognitive distance and reduces the phonetic similarity between the verbal cue and the target L2 word.
@@ 5.3.2 Verbal Cue
Following the highest perplexity score observed with OGR, which results from the use of unconventional keywords, KSS shows the next highest score. This is most likely due to the incorporation of L2 morphological elements in the keywords, as mentioned earlier. These incorporations introduce irregularities that make the model harder to predict, resulting in higher perplexity scores.
@@ 5.4.1 Prompts
Table 2 shows a result on ablating prompt for generating verbal cues. As discussed in Section 3.2, OGR uses two-step approach of generating a story then summarizing while PHT generates verbal cue right away. The result shows that PHT achieves better performance on both metrics, indicating that the two-step approach might be benifical for English-only or verbal cue generations that require multiple keywords. However, in our cross-lingual setting, where there are only two keywords, generating verbal cue right away generates a better verbal cue.
@@ 6.4.1 Correctness
We assess the correctness of recognition and generation response using LLM-as-judge (GPT-4o) Chiang and yi Lee (2023). Previously, Savva et al. (2014) employ Levenshtein distance for assessing correctness. However, as the responses from recognition might involve synonym usage, minor part-of-speech variations, or unintentional typos, relying solely on surface-level string similarity metrics like Levenshtein distance may lead to misleading evaluations. Therefore, we adopt a more semantically aware approach by leveraging GPT-4o as a judge to assess the alignment between the model output and the answer (See Appendix D.4.1 for details).
@@ 6.5.2 Preference Ratings
PHT receives higher ratings than OGR for helpfulness and imageability, while OGR was rated higher for coherence. However, these differences were not statistically significant. Notably, although PHT achieves significantly higher correctness during the generation phase, this does not correlate to helpfulness - consistent with prior work showing subjective preference does not always correlate with verbal cue effectiveness Balepur et al. (2024). In terms of coherence, OGR receives higher ratings because its cue generation transforms a meaningless keyword into a meaningful word, as shown in high modification rate, providing greater flexibility and resulting in more logically coherent cues.
@@ 6.5.3 Case Study
We qualitatively analyze cases where PHT achieves a higher score than KSS and vice versa. The cases where PHT scores higher often involve words containing the consonant /r/, such as reckon and render. PHT uses the first keyword starting with phoneme /l/, which is phonetically close to /r/. However, KSS uses initial keyword starting with /n/, which is more distant in terms of onset alignment. The same pattern occurs with the second keyword.
KSS tends to score higher in cases where lexical flexibility is required, such as with the words felon and mayhem. KSS uses the first keyword “to beat” in its conjugated verb form, while PHT uses the first keyword Peleus, a Greek hero. Notably, OGR also uses the same keyword as KSS, highlighting the limitations of a fixed dictionary compared to the flexibility of human or LLM-generated creativity. These findings suggest that future improvements involve increasing the variety of words in the dictionary pool. (See Appendix D.5 for details).
@@ 7 Conclusion and Future Works
Future research opportunities include: (1) incorporating flexible part-of-speech handling in our keyword sequence retrieval mechanism to improve phonological matches; (2) extending our system to diverse language pairs with varying structural characteristics; and (3) evaluating long-term retention effects in longitudinal studies. These developments will advance mnemonic generation systems for language learners across multiple language families.
@@
Our investigation exhibits four primary constraints. First, we limit our research scope to English-Korean language pairs, necessitating future adaptations for other language combinations with distinct phonological structures and orthographic systems. The cross-lingual IPA transliteration module requires language-specific modifications to accommodate diverse phoneme inventories and phonetic constraints. Second, our evaluation methodology assesses only short-term recall performance rather than longitudinal retention. Future research requires delayed post-tests to evaluate long-term memory consolidation and mnemonic durability. Third, our vocabulary selection derives from standardized test materials targeting advanced-level English learners, potentially limiting PhoniTale’s applicability for beginning learners acquiring common vocabulary. The system’s efficacy may vary with lexical complexity and syllabic structure. Fourth, our Korean dictionary dataset contains only base forms of verbs without their various conjugated forms such as past tense, future tense, or other inflections. This limitation reduces the lexical diversity of potential keywords and constrains our ability to generate more natural-sounding verbal cues.

### v3

@@ Abstract
Vocabulary acquisition poses a significant challenge for second-language (L2) learners, especially when learning typologically distant languages such as English and Korean, where phonological and structural mismatches complicate vocabulary learning. Recently, large language models (LLMs) have been used to generate keyword mnemonics by leveraging similar keywords from a learner’s first language (L1) to aid in acquiring L2 vocabulary. However, most methods still rely on direct IPA-based phonetic matching or employ LLMs without phonological guidance. In this paper, we present PhoniTale, a novel cross-lingual mnemonic generation system that performs IPA-based phonological adaptation and syllable-aware alignment to retrieve L1 keyword sequence and uses LLMs to generate verbal cues. We evaluate PhoniTale through automated metrics and a short-term recall test with human participants, comparing its output to human-written and prior automated mnemonics. Our findings show that PhoniTale consistently outperforms previous automated approaches and achieves quality comparable to human-written mnemonics.
@@ 1 Introduction
Typically, the L1 phrase corresponding to a given L2 term to be memorized is manually designed; however, this is a laborious process that scales poorly, necessitating the development of automated mechanisms to compose these phrases. Methods for automated generation of such keywords began with TransPhoner, which leverages the International Phonetic Alphabet (IPA) International Phonetic Association (1949) and hand-crafted heuristics to retrieve pronunciation-similar L1 keyword for target English words, resulting in significant recall gains Savva et al. (2014).
Leveraging these methods, recent studies employ large language models (LLMs) for automated keyword generation. SmartPhone fed the TransPhoner keyword into GPT-3 to automatically generate verbal cues and DALL·E to generate visual cues Lee and Lan (2023). Lee et al. (2024) introduced an overgenerate-and-rank approach, where LLMs overgenerate keyword sequences and verbal cues, and then rank them according to multiple different criteria. Balepur et al. (2024) aligned mnemonics with user preferences by fine-tuning Llama 2 for personalization and cost-efficiency. Lee et al. (2025) learned latent user and Kanji (Chinese characters in Japanese) traits from a crowd-sourced platform for learning Kanji, and extract rules for constructing mnemonics using an Expectation-Maximization style algorithm.
[…]
@@
In this paper, we introduce a mnemonic generation system for language learning, PhoniTale. Our approach employs a greedy search through phonetically and syllabically approximated L2 sequences to identify the most suitable L1 keyword sequence. Specifically, we first transliterate L2 phonemes into L1-adapted sequences, segment these into syllables according to L1 phonological constraints, and then select keywords that maximize phonetic similarity while preserving syllabic structure. Unlike previous approaches that rely heavily on LLMs for keyword generation, we utilize LLMs only for verbal cue generation, while our specialized modules handle the cross-lingual phonological alignment. This design addresses the unique challenges posed by typologically distant languages, improves scalability, and mitigates hallucination risk. Through systematic evaluation including both automated metrics and human studies with short-term recall tests, we demonstrate that PhoniTale achieves comparable performance to human-authored mnemonics.
@@ 2 Problem Statement
PhoniTale performs the task of retrieving cross-lingual phonologically similar keyword sequence and using them to construct a L1 verbal cue for a given L2 target word, following the process illustrated in Figure 1. Let be a word in the L2, and let denote its meaning. The goal is to retrieve a L1 keyword sequence that are phonologically similar to segments of , and to use to construct a verbal cue , where is the space of natural-language expressions in L1.
@@ 3.1.2 Segmentation
[…]
@@ 3.2 Verbal Cue Generation
Second, we discard the Age-of-Acquisition (AoA) ranking criterion from Lee et al. (2024) as it does not generalize effectively to cross-lingual contexts. The AoA of a word in L2 fails to reliably reflect its familiarity in L1. We retain only the context completeness criterion, calculating this by masking the target word in the verbal cue and prompting GPT-4o OpenAI (2024) to generate five probable candidates. We then compute the average cosine similarity between FastText Bojanowski et al. (2016) embeddings of these candidates and the target word, trained on Korean corpus data Lee (2020). This approach quantifies how effectively the verbal cue provides context for learning the target word’s meaning.
@@ 4.1 Dataset
Keyword Pool
We construct a keyword candidate pool by filtering out non-lexical items such as grammatical particles, suffixes, and sentence-final endings from the Basic Korean Dictionary dataset Ha (2023).
The keyword pool consists of 55,316 unique entries that are phonologically representative and semantically well-formed.
Train
We construct a training dataset of 2,870 English–Korean word pairs with aligned IPA transcriptions. English vocabulary items originate from standardized GRE preparation materials, including official Educational Testing Service guides and commercial resources Magoosh (2021); Princeton Review (2020). Each entry in our dataset includes an English word, its Korean transliteration, and IPA transcriptions for both languages, with syllable-level boundaries annotated in the Korean IPA.
For English, we obtain IPA transcriptions directly from the Oxford University Press (n.d.), which provides standardized phonetic representations widely used in linguistic research.
For Korean, we first extract transliterations of English words from the Aha Dictionary (n.d.). These transliterations are segmented into syllable blocks of the Korean writing system (Hangul), each composed of an initial consonant, a medial vowel, and an optional final consonant. Hangul is often characterized as an alphabetic syllabary, where individual graphemes (jamo) form syllabic blocks (kulja) corresponding to single phonological units Pastore (2019). Each syllable block is then converted into its phonetic representation using the rule-based Hangul-to-IPA conversion method Nam (2022), which encodes standard Korean phonological processes Shin et al. (2012). To identify syllable boundaries within the resulting IPA sequence, we extract the final IPA symbol from each block and annotate it with a binary indicator denoting syllable-final positions. This procedure enables consistent segmentation and alignment of syllable-level IPA representations across English and Korean.
Test
We use the book KSS Gyeong (2020) as our baseline for human-authored verbal cues, designed for native Korean speakers learning English. The vocabulary targets advanced-level standardized tests, including government employee entrance exams, university transfer admissions, the TOEFL Educational Testing Service , and the TEPS Seoul National University . From this vocabulary, we construct a test set of 36 words.
@@ 4.2 Cross-lingual IPA Transliteration
To understand how our model addresses phonological divergence between English and Korean, we analyze attention patterns (See Appendix Figure 6). The visualizations reveal the model’s strategies for cross-linguistic challenges: English affricates decompose into multiple Korean consonants; compatible sounds maintain one-to-one mappings; English diphthongs expand to accommodate Korean’s vowel inventory; and syllable structures adapt to Korean phonotactic constraints. These patterns confirm the model’s ability to dynamically adjust its mapping strategy based on input characteristics.
@@ 4.3 Segmentation
Our model achieves a perfect boundary-level F1 score of 1.00 compared to reference boundaries on system transliteration outputs. This result indicates exceptional precision in identifying syllable junctures within the predicted L1-adapted sequence.
@@ 5.1 Dataset
We use same dataset mentioned in 4.1 for evaluation. We replicate Lee et al. (2024) in a same way for generating cross-lingual verbal cues, except that we translate prompts that were originally written in English with in-context examples for English-to-English learning into Korean, and use in-context examples from KSS to compare with PhoniTale. Hereafter, we refer to the human-authored book as KSS, Lee et al. (2024) as OGR, and PhoniTale as PHT. We use GPT-4o (temperature = 0.7) for both OGR and PHT throughout the entire pipeline to ensure fair comparison, following the temperature setting from Lee et al. (2024).
@@ 5.2.1 Keyword Sequence
We evaluate quality on three aspects: phonetic similarity, keyword omission, and keyword modification. We evaluate phonetic similarity using our IPA-based contrastive model, which measures how closely the concatenated Korean keyword sequence resembles the phonetic form of the English word.
@@ 5.3.1 Keyword Sequence
KSS, authored by human, one possible reason for the low phonetic similarity is the its substitution of L1 meanings for L2 prefixes (e.g., re-, in-) and L2 suffixes (e.g., -cracy). For example, re- is mapped to the L1 word meaning again, and in- is mapped to the L1 word meaning inside.
@@ 5.3.2 Verbal Cue
Following the highest perplexity score observed with OGR, which results from the use of unconventional keywords, KSS shows the next highest score. This is most likely due to the incorporation of L2 morphological elements in the keywords, as mentioned earlier. These incorporation introduce irregularities that make the model harder to predict, resulting in higher perplexity scores. Beyond their surface inclusion, KSS often requires learners to disambiguate polysemous morphemes such as re-, which can mean either again or back depending on the context. These inconsistencies in semantic interpretation and structural mapping increase irregularity in surface realizations, thereby hindering accurate verbal cue generation.
@@ 5.4.1 Prompts
Table 2 shows a result on ablating prompt for generating verbal cues. As discussed in Section 3.2, OGR uses two-step approach of generating a story then summarizing while PHT generates verbal cue right away. The result shows that PHT achieves better performance on both metrics, indicating that the two-step approach might be beneficial for English-only or verbal cue generations that require multiple keywords. However, in our cross-lingual setting, where there are only two keywords, generating verbal cue right away generates a better verbal cue.
@@ 6.4.1 Correctness
We assess the correctness of recognition and generation response using LLM-as-a-judge (GPT-4o) Chiang and yi Lee (2023). Previously, Savva et al. (2014) employ Levenshtein distance for assessing correctness. However, as the responses from recognition might involve synonym usage, minor part-of-speech variations, or unintentional typos, relying solely on surface-level string similarity metrics like Levenshtein distance may lead to misleading evaluations. Therefore, we adopt a more semantically aware approach by leveraging GPT-4o as a judge to assess the alignment between the model output and the answer (See Appendix D.4.1 for details).
@@ 6.5.2 Preference Ratings
PHT receives higher ratings than OGR for helpfulness and imageability, while OGR is rated higher for coherence. However, these differences were not statistically significant. Notably, although PHT achieves significantly higher correctness during generation, this does not correlate with helpfulness. This finding is consistent with prior work showing that subjective preference does not always align with verbal cue effectiveness Balepur et al. (2024). In terms of coherence, OGR receives higher ratings because its cue generation transforms a meaningless keyword into a meaningful word, as shown in high modification rate, providing greater flexibility and resulting in more logically coherent cues.
@@ 6.5.3 Case Study
[…]
However, KSS achieves higher correctness than PHT when keyword sequences are culturally rooted. For example, for felon, KSS adapts the idiom “to administer cudgel strokes” into the verbal cue one who will cudgel-beat, therefore, a felon. This construction is grammatically incorrect because it describes the one doing the beating rather than the one being beaten, yet learners readily reinterpret it as referring to the person who deserves punishment. The effectiveness of this vivid and culturally familiar cue is shown from its higher preference score compared to PHT. In contrast, PHT selects Peleus, a mythological name that preserves phonological alignment but lacks cultural resonance, making it harder to remember. This case illustrates how KSS benefits from culturally rooted and expressive forms, while current LLMs, constrained by grammaticality, struggle to produce such non-standard yet pedagogically effective cues.
@@ 7 Conclusion and Future Works
[…]
Second, we plan to extend the phoneme‐anchored retrieval system to code‐switched speech recognition. In such settings, phonological cues often transcend language boundaries, complicating identification of transliterated loanwords and domain-specific terms. By leveraging IPA-based representations, our approach offers a language-agnostic substrate for capturing cross-lingual phonetic similarity. This can improve recognition of borrowed or specialized vocabulary that deviates from canonical pronunciations, thereby reducing Word Error Rate (WER) in code-switched ASR scenarios. We plan to evaluate this by aligning phonetic units across typologically distant languages and assessing recognition gains for foreign-sounding or morphologically irregular tokens.
Together, these future directions aim to evolve PhoniTale into a more versatile, language‐agnostic tool with applications spanning multilingual mnemonic generation, resource creation, and speech recognition in phonologically diverse or code‐mixed environments.
@@
Our investigation exhibits four primary constraints. First, we limit our research scope to English-Korean language pairs due to the limited availability of training data, necessitating future adaptations for other language combinations with distinct phonological structures and orthographic systems. Second, our evaluation methodology assesses only short-term recall performance rather than longitudinal retention. Future research requires delayed post-tests to evaluate long-term memory consolidation and mnemonic durability. Third, our vocabulary selection derives from standardized test materials targeting advanced-level English learners, potentially limiting PhoniTale’s applicability for beginning learners acquiring common vocabulary. Fourth, while our Korean dictionary dataset includes some conjugated forms, its lexical coverage remains limited. The absence of commonly used loanwords, neologisms, and other everyday variants reduces the pool of potential keywords and constrains the naturalness of generated verbal cues.
Acknowledgments
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP)-Global Data-X Leader HRD program grant funded by the Korea government (MSIT) (IITP-RS-2024-00440626)
@@ B.4 PhoniTale Prompt
The prompt was originally designed in Korean. For reproducibility, we provide both the original and its English translation.
@@
| Game description: In the Story-Chaining Game, players receive a target word candidate set and a keyword set. Their task is to craft a short, concise, single-sentence story that cleverly incorporates these words. The ultimate challenge is to construct a sentence that includes at least one of the target word candidates of your choice and strictly uses the provided keywords in the given order. | |
@@
| 1. Each player receives a set of target word candidates and keywords. | |
@@
| 2. The target word must be chosen first. If there is only one candidate, that word becomes the target word. If multiple candidates are provided, the player must select one. | |
@@
| 3. The target word and keywords must be used to create a short, single-sentence story. | |
@@
| 4. Keywords must appear exactly in the specified order. | |
@@
| 5. The story must include the target word exactly once, and it should be highlighted using angle brackets (< >). | |
@@
| 6. The entire output must be returned in JSON format. | |
@@
| The following are examples of the expected input and output format: | |
@@
| "story": "While peeing in the bush, water <splashed>." | |
@@
| "story": "The mermaid fumbled around and moved indolently." | |